# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 9.3. The proof
We consider each case in turn. Assume first $S=\mathrm{PSL}^{\varepsilon}_{5}(r)$ with $\varepsilon\in\{+,-\}$, so $q\geq r$. For $n=10$ or $15$, up to quasi-equivalence $V$ is the alternating square or the symmetric square of the natural 5-dimensional module, so by Lemma 9.7 we may assume $\varepsilon=$-. For $n=10$, we choose $g$ of order $(r^{5}+1)/(r+1)$, and we calculate directly that $g$ has distinct eigenvalues on $V$. Assume now $n=15$, so by what just proved $g$ is regular on a $10$-dimensional subspace. Since $|N_{L}(\langle g\rangle)/C_{L}(g)|=5$, it follows that there are at most $5$ eigenspaces of dimension $2$, so $d\leq 25$, which implies $|C_{G}(g)|<|G:M|^{1/3}$. (Here we used Lemma 7.12(v) and Lemma 7.1. In the course of the proof, we will often apply these results with no explicit mention.) Assume now $n=23,24$, so we have $|M|<q^{n+3}$. Choosing the original element $g$ and replacing $C=2n+5$ by $C=n+3$ in the above calculation we see that the result follows. (In many cases, below, we will abuse notation and write that $V$ is a certain module, when we really mean that $V$ is quasi-equivalent to that module. This will happen for example in the next paragraph.)
Assume next $S=\mathrm{PSL}^{\varepsilon}_{6}(r)$, so $q\geq r$. In the cases $n=15,21$, by Lemma 9.7 we may assume $\varepsilon=$-. We choose $g$ of order $(r^{5}+1)/(r+1)$. If $n=15$, it follows readily by what proved in the previous paragraph for $\mathrm{PSU}_{5}(r)$ (by restricting to a $10$-dimensional submodule) that $g$ has distinct eigenvalues. If $n=21$, then it similarly follows that $g$ is regular on an $11$-dimensional subspace, so $g$ has at most $5$ eigenspaces of dimension $2$ and $d\leq 31$, which easily gives $|C_{G}(g)|<|G:M|^{1/3}$. For $n=20$, $V$ is the third alternating power. We have $a\leq\binom{6}{2}/3=5$; since $|N_{L}(\langle g\rangle)/C_{L}(g)|\geq 5$, each nontrivial eigenspace has dimension at most $20/5=4$ and this is sufficient.
Assume now $S=\mathrm{PSL}^{\varepsilon}_{7}(r)$ with $n=21$. Since $|N_{L}(\langle g\rangle)/C_{L}(g)|=7$, each eigenspace for $g$ has dimension at most $3$ and $d\leq 63$ and this gives the result.
Assume now $S=\mathrm{PSp}_{4}(r)$ with $n=16$. Then $a\leq 4$ and so each eigenspace has dimension at most $4$, so $d=\dim(C_{\mathrm{GL}_{n}(K)}(g))\leq 64$ and this is sufficient.
@@
Assume now $S=\mathrm{PSp}_{6}(r)$ with $n>8$. Note that $|N_{L}(\langle g\rangle)/C_{L}(g)|=6$. In particular, since $a=\dim(C_{V}(g))\leq n/3$, for $n=13$ we deduce by Lemma 7.1 that $a=1$ and the nontrivial eigenspaces have dimension at most $2$, which gives $|C_{G}(g)|<|G:M|^{1/3}$. By the same argument, for $n=14$ (resp. $n=21$) we see that all eigenspaces of $g$ have dimension at most $2$ (resp. at most $3$), and the result follows.
@@
Assume now $S=G_{2}(r)$. If $n=6$ then $q=r$ is even, $G=\mathrm{Sp}_{6}(q)$ and $S=M$. An element $g\in M$ of order $(q^{3}+1)/(q+1)$ acts irreducibly on $V$, and $g^{G}\cap M=g^{M}$, so $\mathrm{fix}(g,G/M)=|C_{G}(g)|/|C_{M}(g)|=q+1<|G:M|^{1/3}$. The case $n=7$ is entirely analogous; in this case $G=\Omega_{7}(q)$ and $g$ acts irreducibly on a hyperplane. Assume finally $n=14$; since $|N_{G}(\langle g\rangle)|/|C_{G}(g)|=6$ and $a\leq n/3$, we deduce by Lemma 7.1 that all eigenspaces of $g$ have dimension at most $2$, which gives the result.
@@
Assume now $S=\mathrm{PSp}_{4}(7)$ with $n=24,25$; by [32] we have $p\neq 7$ and $G=\mathrm{SL}^{\pm}_{24}(q)$. Let $g\in S$ be of order $7$ and with centralizer of order $4116$; there is only one such $S$-class and so $|g^{S}\cap\langle g\rangle|=6$. We check with GAP that $S$ is generated by two conjugates of $g$. If $n=24$, then $a=\dim(C_{V}(x))\leq n/2=12$ and by Lemmas 7.1 and 7.6 we have that $d=\dim(C_{\mathrm{GL}_{n}(K)}(g))\leq 168$, from which we see that $|C_{G}(x)|<|G:M|^{1/3}$. If $n=25$, then in fact we must have $a\leq 7$, so $d\leq 103$ and this is sufficient.
@@
Assume finally $S=Co_{1}$ with $n=24$. Since $|N_{L}(\langle g\rangle)/C_{L}(g)|=11$ and $a\leq n/2=12$, it follows that all eigenspaces of $g$ have dimension at most $2$, and this gives $|C_{G}(g)|^{3}<|G:M|$.
@@
(Note that the case $S=E_{7}(r)$ with $n=56$ is excluded in view of the assumption $|M|\leq q^{2n+4}$.) Assume first $S=\mathrm{PSU}_{6}(r)$, so $q\geq r$; we have $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=5$. If $n\geq 34$ then $|M|<q^{n+5}$; in the above calculation we may replace $C=2n+5$ by $n+5$ and $B=4$ by $5$ and we see that $|C_{G}(g)|<|G:M|^{1/3}$. The cases $n=15,20,21$ were addressed already in Case 1.
Assume now $S=\mathrm{P}\Omega_{7}(r)$, so $q=r$ and $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=6$. If $n\geq 26$ then $|M|<q^{n-2}$; the usual calculation with $c=1/2$, $B=6$ and C=n2C=n-2 gives the result. If $n=8$ then $G=\Omega_{8}^{+}(q)$ and $S$ is conjugate in $\mathrm{Aut}(G/Z(G))$ to a reducible subgroup, so we can exclude this case. If $n=21$ then $V$ is the alternating square; we have $a=\dim(C_{V}(g))=3$, and so every eigenspace has dimension at most $3$, which implies $|C_{G}(g)|<|G:M|^{1/3}$.
Assume now $S=\mathrm{P}\Omega^{+}_{12}(r)$, so $q=r$. We have $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=6$; if $n\geq 64$ we have $|M|<q^{n+4}$ and the result follows with the usual calculation. If $n=32$ then $V$ is the spin module. The restriction of $V$ to $\text{Spin}^{+}_{8}(q)$ is the sum of four half-spins (two of each type). Letting $g$ be an element of order $q^{4}-1$, it follows that each eigenspace of $g$ on $V$ has dimension at most $4$, which is sufficient.
Assume now $S=\mathrm{P}\Omega^{+}_{8}(r)$, so $q=r$. We have $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=6$. If $n\geq 48$ then $|M|<q^{n}$; the usual calculation with $B=6$ and $C=n$ gives the result. The remaining cases are $n=26,28,35$. If $n=26$ then $p=2$ and $V$ is a composition factor of the alternating square. We have $a=2$, so each nontrivial eigenspace has dimension at most $4$ and this is sufficient. If $n=28$, then $V$ is the alternating square, so $a=4$ and each nontrivial eigenspace has dimension at most $4$, and the result follows. If $n=35$ then $V$ is a composition factor of the symmetric square. We have $a=5$, so all eigenspaces have dimension at most $5$, which gives the result.
@@
If $S=F_{4}(2)$, then $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=12$, so in fact all eigenspaces have dimension at most $2$ and the result follows.
@@
Assume then $S=\mathrm{PSL}^{\pm}_{4}(r)$, so $q\geq r$. If $n\geq 32$, note first that by [61], either $n\geq 44$ or $G=\mathrm{SL}_{n}(q)$ or $\mathrm{SU}_{n}(q^{1/2})$. In any case we have $|M|<q^{n/2+2}$, and the conclusion follows by the usual calculation with $B=4$ and $C=n/2+2$. The remaining cases are $n=6,10,14,15,16,19,20$. We address the case $S=\mathrm{PSL}_{4}^{-}(r)=\mathrm{PSU}_{4}(r)$; the case $S=\mathrm{PSL}_{4}(r)$ is nearly identical. If $n=6$ then $M$ is in class $\mathcal{C}_{8}$ in view of $\mathrm{PSU}_{4}(r)\cong\mathrm{P}\Omega^{-}_{6}(r)$. If $n=10$ then $V$ is the symmetric square and $G=\mathrm{SU}_{10}(q^{1/2})$. Then $a=\dim(C_{V}(g))=2$, so every nontrivial eigenspace has dimension at most $2$, so $d=\dim(C_{\mathrm{GL}_{n}(K)}(g))\leq 20$ and this implies $|C_{G}(g)|<|G:M|^{1/3}$. If $n=14$ then $p=2$ and $V$ is a composition factor of the adjoint module. We have $a=2$, so every nontrivial eigenspace has dimension at most $12/4=3$ and this gives the result. If $n=15$ then $p\neq 2$ and $V$ is the adjoint module; we have $a=3$ and every nontrivial eigenspace has dimension at most $3$ and we are done. If $n=16$ then $p=3$ and $V\cong S^{3}(W)/W$ where $W$ is the natural module. One calculates that $a\leq 4$, so $d\leq 64$ (obtained when $a=0$) and the result follows. If $n=19$ then $a\leq 9$; it follows that in fact $a\leq 7$ and all other eigenspaces have dimension at most $3$, so $d\leq 85$ and this is sufficient. If $n=20$ then it must be $a\leq 8$ and $d\leq 100$, so $|C_{G}(g)|<|G:M|^{1/3}$ also in this case.
@@
Assume now $S=\mathrm{PSp}_{4}(r)$, so $q=r$. If $n\geq 35$ then $|M|<q^{n/3+1}$; we conclude with the usual calculation with $B=4$ and $C=n/3+2$. The remaining cases are $n=10,12,13,14,16,20,24,25,30$. For $i=4,5$, let $W_{i}$ be the natural $i$ dimensional module for $\mathrm{PSp}_{4}(r)$ (recall $\mathrm{PSp}_{4}(r)\cong\Omega_{5}(r)$). If $n=10$ then $V=S^{2}(W_{4})$. We have $a=\dim(C_{V}(g))=2$ and so each eigenspace has dimension at most $2$, which gives $|C_{G}(g)|<|G:M|^{1/3}$. If $n=12$ then $p=5$ and the highest dominant weight is $(1,1)$; it follows that $V$ is a composition factor of $U:=W_{4}\otimes W_{5}$. Let $g\in L$ be an element of order $13$; then $g$ has eigenvalues $x,x^{5},x^{-1},x^{-5}$ on $W_{4}$ and eigenvalues $x^{3},x^{2},x^{-3},x^{-2}$ on $W_{5}$. It follows from an easy calculation that $C_{U}(g)=0$ and each nontrivial eigenspace on $U$ has dimension at most $2$. This implies that $d\leq 24$, which gives $|C_{G}(g)|<|G:M|^{1/3}$. If $n=13$ (resp. $n=14$) then $p=5$ (resp. $p\neq 5$) and $V$ is a composition factor of $S^{2}(W_{5})$. We have $a=1$ (resp. $a=2$) and each nontrivial eigenspace has dimension at most $12/4=3$, which is sufficient. If $n=16$ then the highest dominant weight is $(1,1)$; as in the case $n=12$ we have that $V$ is a composition factor of $U:=W_{4}\otimes W_{5}$. We easily calculate that $C_{U}(g)=0$, and in particular each nontrivial eigenspace on $U$ has dimension at most $4$; it follows that $d\leq 64$, which gives the result. If $n=20$ then as in the $\mathrm{PSL}^{\pm}_{4}(r)$-case we see that $d\leq 100$ and the result follows. If $n=25$ then $a\leq 12$, so in fact $a\leq 9$ and so $d\leq 145$ and we conclude. If $n=24$ then $p=7$ and the highest dominant weight is $(1,2)$ (where the second weight corresponds to the short root); in particular $W$ is a composition factor of $U:=W_{4}\otimes A$ where $A$ is a $14$-dimensional composition factor of $S^{2}(W_{5})$. Let $g$ be an element of order $25$. We calculate that $\dim(C_{A}(g))=2$ and each nontrivial eigenspace has dimension $1$; it follows immediately that each eigenspace on $U$ has dimension at most $4$, from which $d\leq 96$ and $|C_{G}(g)|<|G:M|^{1/3}$.
@@
If $S\neq\mathrm{PSL}^{\pm}_{4}(r),\mathrm{PSp}_{4}(r)$ then $n\geq 25$ and $V=W\otimes W^{(q)}$ where $W$ is the natural module for $S$. Since $g$ has distinct eigenvalues on $W$, it follows that each eigenspace of $g$ on $V$ has dimension at most $\dim(W)=n^{1/2}$, which is sufficient. Assume now $S=\mathrm{PSL}^{\pm}_{4}(r)$. If $n=16$, we see that $C_{V}(g)=0$ and we see that each eigenspace has dimension at most $3$ (in order to prove this, it is enough to check that at least one eigenspace has dimension at most $3$, since the eigenspaces are permuted in orbits of size $4$). This gives the result. If $n=64$ then we see similarly that each eigenspace of $g$ on $W\otimes W^{(q)}$ has dimension at most $3$, from which each eigenspace on $V=W\otimes W^{(q)}\otimes W^{(q^{2})}$ has dimension at most $12$ and this is enough. The case $S=\mathrm{PSp}_{4}(r)$ can be handled by the same argument.
@@
By [32], we see that for $n\leq 27$, either the modular character tables of $S$ and its covers are available in [8], or $S=\mathrm{PSp}_{4}(7)$ (with $n=24,25$) or $S=Co_{1}$ (with $n=24$). If $S=\mathrm{PSp}_{4}(7)$ then $G=\mathrm{SL}^{\pm}_{n}(q)$, in which case we already have the result for $n\geq 22$. If $S=Co_{1}$, we have $p=q=23$ and $M=2.Co_{1}$ and $G=\Omega^{\varepsilon}_{n}(q)$. In this case it is convenient to change our choice of the element $g$. The restriction of $V$ to $H=Co_{2}$ has composition factors of degree $23$ and $1$. Now the modular character table of $Co_{2}$ is available in [8]; we see that an element $g$ of order $11$ has $10$ nontrivial eigenspaces of dimension $2$ and fixed space of dimension $4$. Therefore $d=56$ and $d^{\prime}=d/2-4/2=26$, where $d^{\prime}$ is the dimension of the centralizer of $g$ in $\mathrm{SO}_{n}(K)$, which implies $|M||C_{G}(g)|^{3}<|G|$.
@@
Assume first $f\leq n/10$, so $d\leq 0.272n^{2}$. Moreover, letting $R$ denote the number of Jordan blocks of $g$, we have $R\leq 3f+(n-3f)/5\leq 11n/25$. Since $f\geq 4$, we deduce that $n\geq 10f\geq 40$. Using $|M|\leq q^{2n+5}$ and Lemmas 7.10 and 7.11 we calculate similarly to above that $|M||C_{G}(g)|^{3}<|G|$, unless $G=\Omega^{\varepsilon}_{n}(q)$ with $40\leq n\leq 43$. In these exceptional cases, since $f\leq n/10$ we have $f=4$. We have $\dim(V_{0})\leq 12$ and all Jordan blocks on $U$ have odd size $|g|\geq 5$. If the Jordan block on $V_{0}$ have size at least $2$, it is easy to calculate that $d\leq n^{2}/5+96$. Using $R\leq 11n/25$, we see that $|M||C_{G}(g)|^{3}<|G|$ for $n\geq 39$. The remaining case is where $g$ is trivial on $V_{0}$. Then by [56, Theorem 3.1] we have $d^{\prime}=d/2-R/2$, where $d^{\prime}$ is the dimension of the centralizer of $g$ in $\mathrm{SO}_{n}(K)$, so $d^{\prime}+R=d/2+R/2$. Using as above $R\leq 22n/50$ and $d\leq 0.272n^{2}$, we see that $|M||C_{G}(g)|^{3}<|G|$ for $n\geq 36$, and so this case is done.

### v2

@@ 9.3. The proof
We consider each case in turn. Assume first $S=\mathrm{PSL}^{\varepsilon}_{5}(r)$ with $\varepsilon\in\{+,-\}$, so $q\geq r$. For $n=10$ or $15$, up to quasi-equivalence $V$ is the alternating square or the symmetric square of the natural 5-dimensional module, so by Lemma 9.7 we may assume $\varepsilon=$-. For $n=10$, we choose $g$ of order $(r^{5}+1)/(r+1)$, and we calculate directly that $g$ has distinct eigenvalues on $V$. Assume now $n=15$, so by what just proved $g$ is regular on a $10$-dimensional subspace. Since $|N_{L}(\langle g\rangle)/C_{L}(g)|=5$, it follows that there are at most 55 eigenspaces of dimension 22, so $d\leq 25$, which implies $|C_{G}(g)|<|G:M|^{1/3}$. (Here we used Lemma 7.12(v) and Lemma 7.1. In the course of the proof, we will often apply these results with no explicit mention.) Assume now $n=23,24$, so we have $|M|<q^{n+3}$. Choosing the original element $g$ and replacing $C=2n+5$ by $C=n+3$ in the above calculation we see that the result follows. (In many cases, below, we will abuse notation and write that $V$ is a certain module, when we really mean that $V$ is quasi-equivalent to that module. This will happen for example in the next paragraph.)
Assume next $S=\mathrm{PSL}^{\varepsilon}_{6}(r)$, so $q\geq r$. In the cases $n=15,21$, by Lemma 9.7 we may assume $\varepsilon=$-. We choose $g$ of order $(r^{5}+1)/(r+1)$. If $n=15$, it follows readily by what proved in the previous paragraph for $\mathrm{PSU}_{5}(r)$ (by restricting to a $10$-dimensional submodule) that $g$ has distinct eigenvalues. If $n=21$, then it similarly follows that $g$ is regular on an $11$-dimensional subspace, so $g$ has at most 55 eigenspaces of dimension 22 and $d\leq 31$, which easily gives $|C_{G}(g)|<|G:M|^{1/3}$. For $n=20$, $V$ is the third alternating power. We have $a\leq\binom{6}{2}/3=5$; since $|N_{L}(\langle g\rangle)/C_{L}(g)|\geq 5$, each nontrivial eigenspace has dimension at most $20/5=4$ and this is sufficient.
Assume now $S=\mathrm{PSL}^{\varepsilon}_{7}(r)$ with $n=21$. Since $|N_{L}(\langle g\rangle)/C_{L}(g)|=7$, each eigenspace for $g$ has dimension at most 33 and $d\leq 63$ and this gives the result.
Assume now $S=\mathrm{PSp}_{4}(r)$ with $n=16$. Then $a\leq 4$ and so each eigenspace has dimension at most 44, so $d=\dim(C_{\mathrm{GL}_{n}(K)}(g))\leq 64$ and this is sufficient.
@@
Assume now $S=\mathrm{PSp}_{6}(r)$ with $n>8$. Note that $|N_{L}(\langle g\rangle)/C_{L}(g)|=6$. In particular, since $a=\dim(C_{V}(g))\leq n/3$, for $n=13$ we deduce by Lemma 7.1 that $a=1$ and the nontrivial eigenspaces have dimension at most 22, which gives $|C_{G}(g)|<|G:M|^{1/3}$. By the same argument, for $n=14$ (resp. $n=21$) we see that all eigenspaces of $g$ have dimension at most 22 (resp. at most 33), and the result follows.
@@
Assume now $S=G_{2}(r)$. If $n=6$ then $q=r$ is even, $G=\mathrm{Sp}_{6}(q)$ and $S=M$. An element $g\in M$ of order $(q^{3}+1)/(q+1)$ acts irreducibly on $V$, and $g^{G}\cap M=g^{M}$, so $\mathrm{fix}(g,G/M)=|C_{G}(g)|/|C_{M}(g)|=q+1<|G:M|^{1/3}$. The case $n=7$ is entirely analogous; in this case $G=\Omega_{7}(q)$ and $g$ acts irreducibly on a hyperplane. Assume finally $n=14$; since $|N_{G}(\langle g\rangle)|/|C_{G}(g)|=6$ and $a\leq n/3$, we deduce by Lemma 7.1 that all eigenspaces of $g$ have dimension at most 22, which gives the result.
@@
Assume now $S=\mathrm{PSp}_{4}(7)$ with $n=24,25$; by [32] we have $p\neq 7$ and $G=\mathrm{SL}^{\pm}_{24}(q)$. Let $g\in S$ be of order 77 and with centralizer of order $4116$; there is only one such $S$-class and so $|g^{S}\cap\langle g\rangle|=6$. We check with GAP that $S$ is generated by two conjugates of $g$. If $n=24$, then $a=\dim(C_{V}(x))\leq n/2=12$ and by Lemmas 7.1 and 7.6 we have that $d=\dim(C_{\mathrm{GL}_{n}(K)}(g))\leq 168$, from which we see that $|C_{G}(x)|<|G:M|^{1/3}$. If $n=25$, then in fact we must have $a\leq 7$, so $d\leq 103$ and this is sufficient.
@@
Assume finally $S=Co_{1}$ with $n=24$. Since $|N_{L}(\langle g\rangle)/C_{L}(g)|=11$ and $a\leq n/2=12$, it follows that all eigenspaces of $g$ have dimension at most 22, and this gives $|C_{G}(g)|^{3}<|G:M|$.
@@
(Note that the case $S=E_{7}(r)$ with $n=56$ is excluded in view of the assumption $|M|\leq q^{2n+4}$.) Assume first $S=\mathrm{PSU}_{6}(r)$, so $q\geq r$; we have $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=5$. If $n\geq 34$ then $|M|<q^{n+5}$; in the above calculation we may replace $C=2n+5$ by $n+5$ and $B=4$ by 55 and we see that $|C_{G}(g)|<|G:M|^{1/3}$. The cases $n=15,20,21$ were addressed already in Case 1.
Assume now $S=\mathrm{P}\Omega_{7}(r)$, so $q=r$ and $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=6$. If $n\geq 26$ then $|M|<q^{n-2}$; the usual calculation with $c=1/2$, $B=6$ and C=n2C=n-2 gives the result. If $n=8$ then $G=\Omega_{8}^{+}(q)$ and $S$ is conjugate in $\mathrm{Aut}(G/Z(G))$ to a reducible subgroup, so we can exclude this case. If $n=21$ then $V$ is the alternating square; we have $a=\dim(C_{V}(g))=3$, and so every eigenspace has dimension at most 33, which implies $|C_{G}(g)|<|G:M|^{1/3}$.
Assume now $S=\mathrm{P}\Omega^{+}_{12}(r)$, so $q=r$. We have $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=6$; if $n\geq 64$ we have $|M|<q^{n+4}$ and the result follows with the usual calculation. If $n=32$ then $V$ is the spin module. The restriction of $V$ to $\text{Spin}^{+}_{8}(q)$ is the sum of four half-spins (two of each type). Letting $g$ be an element of order $q^{4}-1$, it follows that each eigenspace of $g$ on $V$ has dimension at most 44, which is sufficient.
Assume now $S=\mathrm{P}\Omega^{+}_{8}(r)$, so $q=r$. We have $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=6$. If $n\geq 48$ then $|M|<q^{n}$; the usual calculation with $B=6$ and $C=n$ gives the result. The remaining cases are $n=26,28,35$. If $n=26$ then $p=2$ and $V$ is a composition factor of the alternating square. We have $a=2$, so each nontrivial eigenspace has dimension at most 44 and this is sufficient. If $n=28$, then $V$ is the alternating square, so $a=4$ and each nontrivial eigenspace has dimension at most 44, and the result follows. If $n=35$ then $V$ is a composition factor of the symmetric square. We have $a=5$, so all eigenspaces have dimension at most 55, which gives the result.
@@
If $S=F_{4}(2)$, then $|N_{L}(\langle g\rangle)|/|C_{L}(g)|=12$, so in fact all eigenspaces have dimension at most 22 and the result follows.
@@
Assume then $S=\mathrm{PSL}^{\pm}_{4}(r)$, so $q\geq r$. If $n\geq 32$, note first that by [61], either $n\geq 44$ or $G=\mathrm{SL}_{n}(q)$ or $\mathrm{SU}_{n}(q^{1/2})$. In any case we have $|M|<q^{n/2+2}$, and the conclusion follows by the usual calculation with $B=4$ and $C=n/2+2$. The remaining cases are $n=6,10,14,15,16,19,20$. We address the case $S=\mathrm{PSL}_{4}^{-}(r)=\mathrm{PSU}_{4}(r)$; the case $S=\mathrm{PSL}_{4}(r)$ is nearly identical. If $n=6$ then $M$ is in class $\mathcal{C}_{8}$ in view of $\mathrm{PSU}_{4}(r)\cong\mathrm{P}\Omega^{-}_{6}(r)$. If $n=10$ then $V$ is the symmetric square and $G=\mathrm{SU}_{10}(q^{1/2})$. Then $a=\dim(C_{V}(g))=2$, so every nontrivial eigenspace has dimension at most 22, so $d=\dim(C_{\mathrm{GL}_{n}(K)}(g))\leq 20$ and this implies $|C_{G}(g)|<|G:M|^{1/3}$. If $n=14$ then $p=2$ and $V$ is a composition factor of the adjoint module. We have $a=2$, so every nontrivial eigenspace has dimension at most $12/4=3$ and this gives the result. If $n=15$ then $p\neq 2$ and $V$ is the adjoint module; we have $a=3$ and every nontrivial eigenspace has dimension at most 33 and we are done. If $n=16$ then $p=3$ and $V\cong S^{3}(W)/W$ where $W$ is the natural module. One calculates that $a\leq 4$, so $d\leq 64$ (obtained when $a=0$) and the result follows. If $n=19$ then $a\leq 9$; it follows that in fact $a\leq 7$ and all other eigenspaces have dimension at most 33, so $d\leq 85$ and this is sufficient. If $n=20$ then it must be $a\leq 8$ and $d\leq 100$, so $|C_{G}(g)|<|G:M|^{1/3}$ also in this case.
@@
Assume now $S=\mathrm{PSp}_{4}(r)$, so $q=r$. If $n\geq 35$ then $|M|<q^{n/3+1}$; we conclude with the usual calculation with $B=4$ and $C=n/3+2$. The remaining cases are $n=10,12,13,14,16,20,24,25,30$. For $i=4,5$, let $W_{i}$ be the natural $i$ dimensional module for $\mathrm{PSp}_{4}(r)$ (recall $\mathrm{PSp}_{4}(r)\cong\Omega_{5}(r)$). If $n=10$ then $V=S^{2}(W_{4})$. We have $a=\dim(C_{V}(g))=2$ and so each eigenspace has dimension at most 22, which gives $|C_{G}(g)|<|G:M|^{1/3}$. If $n=12$ then $p=5$ and the highest dominant weight is $(1,1)$; it follows that $V$ is a composition factor of $U:=W_{4}\otimes W_{5}$. Let $g\in L$ be an element of order $13$; then $g$ has eigenvalues $x,x^{5},x^{-1},x^{-5}$ on $W_{4}$ and eigenvalues $x^{3},x^{2},x^{-3},x^{-2}$ on $W_{5}$. It follows from an easy calculation that $C_{U}(g)=0$ and each nontrivial eigenspace on $U$ has dimension at most 22. This implies that $d\leq 24$, which gives $|C_{G}(g)|<|G:M|^{1/3}$. If $n=13$ (resp. $n=14$) then $p=5$ (resp. $p\neq 5$) and $V$ is a composition factor of $S^{2}(W_{5})$. We have $a=1$ (resp. $a=2$) and each nontrivial eigenspace has dimension at most $12/4=3$, which is sufficient. If $n=16$ then the highest dominant weight is $(1,1)$; as in the case $n=12$ we have that $V$ is a composition factor of $U:=W_{4}\otimes W_{5}$. We easily calculate that $C_{U}(g)=0$, and in particular each nontrivial eigenspace on $U$ has dimension at most 44; it follows that $d\leq 64$, which gives the result. If $n=20$ then as in the $\mathrm{PSL}^{\pm}_{4}(r)$-case we see that $d\leq 100$ and the result follows. If $n=25$ then $a\leq 12$, so in fact $a\leq 9$ and so $d\leq 145$ and we conclude. If $n=24$ then $p=7$ and the highest dominant weight is $(1,2)$ (where the second weight corresponds to the short root); in particular $W$ is a composition factor of $U:=W_{4}\otimes A$ where AA is a $14$-dimensional composition factor of $S^{2}(W_{5})$. Let $g$ be an element of order $25$. We calculate that $\dim(C_{A}(g))=2$ and each nontrivial eigenspace has dimension 11; it follows immediately that each eigenspace on $U$ has dimension at most 44, from which $d\leq 96$ and $|C_{G}(g)|<|G:M|^{1/3}$.
@@
If $S\neq\mathrm{PSL}^{\pm}_{4}(r),\mathrm{PSp}_{4}(r)$ then $n\geq 25$ and $V=W\otimes W^{(q)}$ where $W$ is the natural module for $S$. Since $g$ has distinct eigenvalues on $W$, it follows that each eigenspace of $g$ on $V$ has dimension at most $\dim(W)=n^{1/2}$, which is sufficient. Assume now $S=\mathrm{PSL}^{\pm}_{4}(r)$. If $n=16$, we see that $C_{V}(g)=0$ and we see that each eigenspace has dimension at most 33 (in order to prove this, it is enough to check that at least one eigenspace has dimension at most 33, since the eigenspaces are permuted in orbits of size 44). This gives the result. If $n=64$ then we see similarly that each eigenspace of $g$ on $W\otimes W^{(q)}$ has dimension at most 33, from which each eigenspace on $V=W\otimes W^{(q)}\otimes W^{(q^{2})}$ has dimension at most $12$ and this is enough. The case $S=\mathrm{PSp}_{4}(r)$ can be handled by the same argument.
@@
By [32], we see that for $n\leq 27$, either the modular character tables of $S$ and its covers are available in [8], or $S=\mathrm{PSp}_{4}(7)$ (with $n=24,25$) or $S=Co_{1}$ (with $n=24$). If $S=\mathrm{PSp}_{4}(7)$ then $G=\mathrm{SL}^{\pm}_{n}(q)$, in which case we already have the result for $n\geq 22$. If $S=Co_{1}$, we have $p=q=23$ and $M=2.Co_{1}$ and $G=\Omega^{\varepsilon}_{n}(q)$. In this case it is convenient to change our choice of the element $g$. The restriction of $V$ to $H=Co_{2}$ has composition factors of degree $23$ and 11. Now the modular character table of $Co_{2}$ is available in [8]; we see that an element $g$ of order $11$ has $10$ nontrivial eigenspaces of dimension 22 and fixed space of dimension 44. Therefore $d=56$ and $d^{\prime}=d/2-4/2=26$, where $d^{\prime}$ is the dimension of the centralizer of $g$ in $\mathrm{SO}_{n}(K)$, which implies $|M||C_{G}(g)|^{3}<|G|$.
@@
Assume first $f\leq n/10$, so $d\leq 0.272n^{2}$. Moreover, letting $R$ denote the number of Jordan blocks of $g$, we have $R\leq 3f+(n-3f)/5\leq 11n/25$. Since $f\geq 4$, we deduce that $n\geq 10f\geq 40$. Using $|M|\leq q^{2n+5}$ and Lemmas 7.10 and 7.11 we calculate similarly to above that $|M||C_{G}(g)|^{3}<|G|$, unless $G=\Omega^{\varepsilon}_{n}(q)$ with $40\leq n\leq 43$. In these exceptional cases, since $f\leq n/10$ we have $f=4$. We have $\dim(V_{0})\leq 12$ and all Jordan blocks on $U$ have odd size $|g|\geq 5$. If the Jordan block on $V_{0}$ have size at least 22, it is easy to calculate that $d\leq n^{2}/5+96$. Using $R\leq 11n/25$, we see that $|M||C_{G}(g)|^{3}<|G|$ for $n\geq 39$. The remaining case is where $g$ is trivial on $V_{0}$. Then by [56, Theorem 3.1] we have $d^{\prime}=d/2-R/2$, where $d^{\prime}$ is the dimension of the centralizer of $g$ in $\mathrm{SO}_{n}(K)$, so $d^{\prime}+R=d/2+R/2$. Using as above $R\leq 22n/50$ and $d\leq 0.272n^{2}$, we see that $|M||C_{G}(g)|^{3}<|G|$ for $n\geq 36$, and so this case is done.